Mass Data and Calculations Table Aluminum Brass Iron M1 (kg) | M2(kg) | M3 (kg) | M4 (kg) | M(kg) | ơn--1 (kg) | 때(kg) 0.0863 00864 0.086 0086 0.08631 1.436,10 .18 l0 02699 0.2699 0.26 0.260 0.2696945830 .2810 Zero Reading of the calipers Length Data and Calculations Table 0.0315 0.0314 0035 0.0349 0.00165 8.250 Aluminum 0.0318 10.0318 | 0.0310|0.0316 |0.0318 |1.633x1018-16 SUP 00315 0.0315 lo031S 0,0316 0,0315 ч L0824(1014.12(10 Iron Diameter Data and Calculations Table 10.0318. | О-O315010, 03144 |0,03150|0.03244|0.00165|8. 2 Slo Aluminum Brass 8 Iron
media%2Fcd4%2Fcd4cddd3-fd89-4f19-848d-b3
QUESTIONS 1. Consider the uncertainty in the measured value of p to be given by ap. Taking the one decimal place of αρ as the least significant digit in ρ, how many significant figures measurements of p? are indicated for each of the 2. What is the accuracy of your determination of the density for each metal? State clearly what quantity describes the accuracy of your measurements of the density.Using the tables answer the question 1, (under questions)
Answer #1

1. Consider the uncertainty in the measured value of (\rho) to be given by \alpha\rho.

Taking the one decimal place of \alpha\rho as the least significant digit in \rho, how many significant figures are indicated for each of the measurement of '\rho'.

(a) For Aluminium, we have

\rhoexp = 2764 kg/m3 = 2.764 x 103 kg/m3         ;       Significant figure = 4

\alpha\rho = 167 kg/m3 = 1.67 x 102 kg/m3              ;       Significant figure = 3

\rhoknown = 2700 kg/m3 = 2.70 x 103 kg/m3       ;       Significant figure = 2

(b) For Brass, we have

\rhoexp = 8387 kg/m3 = 8.387 x 103 kg/m3         ;       Significant figure = 4

\alpha\rho = 86 kg/m3 = 8.6 x 101 kg/m3                ;       Significant figure = 2

\rhoknown = 8400 kg/m3 = 8.4 x 103 kg/m3       ;      Significant figure = 2

(c) For Iron, we have

\rhoexp = 7937 kg/m3 = 7.937 x 103 kg/m3       ;        Significant figure = 4

\alpha\rho = 22 kg/m3 = 2.2 x 101 kg/m3                ;        Significant figure = 2

\rhoknown = 7850 kg/m3 = 7.85 x 103 kg/m3       ;        Significant figure = 3